1.an ideal gas engine oprates in a car not cycle between 227° and 127°.it absorbs 2.095*(10)5 J at the higher temperature.what amount of work is the engine theoretically capable of performing?

Grade:11

1 Answers

Shailendra Kumar Sharma
6 years ago
Maximum efficiency of system will be = 1-\frac{Tsink}{Tsource}
=1-(273+127)/(273+227)
=0.2
So the work delivered will be 2.095*10*5*.2 =20.95Joules